OpenStudy (anonymous):

If a pro basketball player has a vertical leap of about 35 inches, what is his hang time? use the hang time as a function V=48T^2

OpenStudy (anonymous):

Is V = vertical leap?

OpenStudy (anonymous):

yes i think

OpenStudy (anonymous):

\[T = \sqrt{V\div48}\]

OpenStudy (anonymous):

is that the answer because i do not know how to do this problem?

OpenStudy (anonymous):

does anybody else know what to do?

OpenStudy (anonymous):

Given that v is the vertical leap and t is the time you're looking for. Sub in v so that... \[t = \sqrt{35/48} = 0.85391256 seconds\]